OpenStudy (anonymous):

Evaluate f(x) = 2 x 4^x for x = -2 a. 0.225 b. 16 c. 0.125 d. -8

OpenStudy (anonymous):

So plugging in the -2 for gives you 2 x 4^(-2) This becomes 2 x 1/4^2 Why is it 1/4^2? Because a negative power means 1 over the value to that power. In this case the value is 4 and the power is 2. Simplify further to get 2 x 1/16= 2/16=1/8=.0125